# SonifyAR: Context-Aware Sound Generation in Augmented Reality

### v1

@@ Abstract.
Sound plays a crucial role in enhancing user experience and immersiveness in Augmented Reality (AR). However, current platforms lack support for AR sound authoring due to limited interaction types, challenges in collecting and specifying context information, and difficulty in acquiring matching sound assets. We present SonifyAR, an LLM-based AR sound authoring system that generates context-aware sound effects for AR experiences. SonifyAR expands the current design space of AR sound and implements a Programming by Demonstration (PbD) pipeline to automatically collect contextual information of AR events, including virtual content semantics and real world context. This context information is then processed by a large language model to acquire sound effects with Recommendation, Retrieval, Generation, and Transfer methods. To evaluate the usability and performance of our system, we conducted a user study with eight participants and created five example applications, including an AR-based science experiment, an improving case for AR headset safety, and an assisting example for low vision AR users.
@@
$c$s: Human-centered computing Interaction designccs: Human-centered computing Interactive systems and tools
Figure 1. SonifyAR is an AR sound authoring pipeline that generates context-matching sounds for AR events in situ. With SonifyAR, the user can first interact with the AR Scene, while SonifyAR identifies potential sound-producing events and collects context information as text. This context text, including the real-world environment information like surface material, is then processed by LLM, which guides the sound acquisition process using multiple models and methods. The acquired sounds are then used to sonify AR events.
@@
SonifyAR overview
@@ 1. Introduction
Sound is a critical but often overlooked element in Augmented Reality (AR). AR-based sound can improve immersion and overall user experience (roginska2017immersive, ) in applications such as depth perception and task completion (zhou_role_2007, ), search and navigation (ruminski2015experimental, ), and even assistance for people with low vision  (ribeiro_auditory_2012, ). Despite its importance, current AR authoring platforms like Reality Composer (realitycomposer, ), Adobe Aero (adobeaero, ), and Unity Mars (UnityMARS2023, ) provide only rudimentary sound support. Specifically, we identified three critical gaps with existing AR authoring tools:
@@
To address these challenges, we present SonifyAR: a context-aware AR sound authoring system using GPT-4 (openai2024gpt4, ) and a text2audio diffusion model called AudioLDM (liu2023audioldm, ). SonifyAR makes the following key technical advancements:
First, context collection using PbD. SonifyAR adopts a Programming by Demonstration (PbD) (10.1145/291080.291104, ) pipeline to simplify the specification of complex AR interactions. The PbD pipeline enables users to demonstrate AR sound interactions while the system automatically detects the action and collects context information. For example, if a creator wants to sonify the stomping of a walking robot, they can position the virtual robot on the target (physical) surface. As the robot walks, the collision between robot’s feet and the surface, as well as the context information like the robot’s attributes and the surface’s material, is captured by SonifyAR .
Second, context as text. To utilize AR context information stemming from multiple sources (e.g. user action, virtual object, real-world environment) and in different formats (e.g. categorical, 3D shape, image), we use text as the universal medium to encompass all context information. By describing all context with text and fitting different parts into a template (”This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities]”), we represent context information of AR interactions as text snippets.
@@
To author AR sound with SonifyAR, the creator initiates AR interactions and the system automatically lists sound options based on context. For example, imagine trying to make a virtual tea cup chime crisply when being slid on a table (Figure 1). In traditional approaches, the creator would need to manually specify this sliding action and find a sound asset to match the chiming ceramic cup. But with SonifyAR, the creator could demonstrate this sliding action and the chime sounds are generated automatically for the creator to choose from.
To explore the potential of SonifyAR, we conducted two evaluations: first, a qualitative user study with eight designers to examine the usability of the SonifyAR prototype; second, we apply SonifyAR across five key user scenarios: from AR education to improving AR headset safety. Our user study highlights usability of the authoring interface and helpfulness of the generated sounds while identifying space of improvements in terms of generated sound quality. Our application scenarios demonstrate the breadth and potential of SonifyAR incorporating automatically acquired sounds seamlessly into various AR experiences—a practice that would otherwise take significant time and effort.
@@ 2.1. AR Sound Authoring
Sound in AR provides many benefits, from directing the user’s attention and enhancing immersion (roginska2017immersive, ; zhou_role_2007, ) to creating interactive time-varying experiences and increasing accessibility (serafin_sonic_2018, ; ruminski2015experimental, ). However, as mentioned, the current landscape of AR authoring tools and platforms reveals a deficiency of support in AR sound authorship.
The existing AR authoring tools show diversity in their authoring capability (monteiro2023teachable, ; nebeling_trouble_2018, ). Some are easy to use, require no coding skills, and can only attach user-provided sound with certain AR event triggers (haloar, ; arvid, ; adobeaero, ; realitycomposer, ). Some are heavily code-based and require significant technical skills, but provide substantial flexibility in designing AR sound interactions, including the ability to set parameters like sound decay and code-specific conditions for activating sound assets(unity-website, ; unreal-engine-website, ). One common design in these tools is a Programming by Specification (PbS) methodology (monteiro2023teachable, ), where creators define AR sound content and play conditions during the design phase. This allows creators to precisely define the conditions for triggers and the consequent sound effects.
As mentioned in the Introduction, we identified critical gaps with existing AR authoring tools. These gaps are partially due to the limitation of the PbS methodology, since the triggers specified in the design stage naturally lack richness in real-world context, which is crucial for AR sound to be realistic. Also, the trade-offs between supported trigger types and the tool’s usability make most tools lack coverage for many sound-producing AR events. Additionally, all sound assets need to be manually selected or created, tasking creators with sourcing suitable sound effects to match the specific triggers. The SonifyAR system aims to address these gaps.
@@ 2.2. Sound Acquisition
Although no work has explored sound generation with AR events as input, sound generation conditioned from other input modalities like images, 3D models, text, and videos has already been widely applied. Various cross-modal generative models like RNN (Zhou_2018_CVPR, ), GAN (engel2019gansynth, ; kumar2019melgan, ; ghose2022foleygan, ), VAE (dhariwal2020jukebox, ), and most recently diffusion models (kong2020diffwave, ; liu2023audioldm, ; yang2023diffsound, ), are utilized to generate sound. Despite different model architecture, one common training goal is the mapping between latent representations of input modalities and output sound, thus sounds can be generated from prompts like images (10096023, ), videos (Zhou_2018_CVPR, ; ghose2022foleygan, ), and text (liu2023audioldm, ; yang2023diffsound, ). For example, FoleyGAN (ghose2022foleygan, ) conditions action sequences of input videos leading to the generation of visually aligned realistic soundtracks. AudioLDM (liu2023audioldm, ) utilize CLAP (wu2023large, ) based latent space to embed input text and generate matching sound with diffusion model. Unlike the cross-modal generative methods, another notable line of work is physics-based sound synthesis (raghuvanshi_interactive_2006, ; roodaki_sonifeye_2017, ; ren_example-guided_2013, ; jin_deep-modal_2020, ; jin_neuralsound_2022, ; liu_sound_2021, ; diaz_rigid-body_2022, ). Traditionally using methods like spring-mass model (raghuvanshi_interactive_2006, ) and recently improved with deep learning (jin_deep-modal_2020, ), these synthesis methods can provide realistic sound simulation given 3D model and material properties.
Besides synthesizing, retrieval methods can also acquire matching sound for input conditions. For example, Oncescu et al. (Oncescu21a, ) trained cross-modal embedding models to support text-to-audio retrieval. Multi-modality models like VAST  (chen2023vast, ) also supports text-to-audio retrieval that could potentially be used to provide sound assets for AR experiences. A recent work called Soundify (lin2023soundify, ) utilizes retrieved sound assets to sonify videos. They claim that although sound retrieval do not ensure coverage for all input descriptions, the retrieved sound assets usually surpass synthesis results in terms of quality. In this case, combining generation and retrieval in the sound acquisition pipeline seems to be a natural method that integrates merits of both. We are going to explore this possibility in this work.
Unlike other existing sound acquisition methods, AR poses unique challenges in the matching of input condition and output sound. The multi-modal nature of AR interactions, including the information of the virtual object (e.g. a toy robot with a walking animation), real environment (e.g. a living room space with a wooden table in the middle), and user intent (e.g. user taps on a virtual model), need to be carefully processed into the sound acquisition pipeline to result in matching sound assets.
@@ 2.3. Context-awareness in AR
In AR, context-awareness is an important property that allows AR content to be dynamically adjusted according to context information, such as location, scene semantics, and human factors. Context-awareness ensures that the application adapts not only to the user but also to their environment and specific needs, making the application more user-friendly and immersive (krings_development_2020, ). Previous research has explored various aspects of context-awareness in AR, highlighting its importance across different applications.
For example, the popular AR game Pokemón Go shows game contents based on user’s physical location. A more common practice in AR research is adapting virtual contents to scene semantics—essentially understanding the objects and their locations in the user’s environment. Qian et al. (qian2022scalar, ) introduced an authoring system that helps user create adaptive AR experiences to surrounding scene semantics. Lang et al. (lang2019virtual, ) analyzes scene semantics to guide the positioning of virtual agents. Similarly, Liang et al. (liang2021scene, ) places virtual pets into AR based on scene geometry and semantics. This realm of similar work (han2020live, ; tahara_retargetable_2020, ) usually captures and analyzes surrounding environment, and designs algorithms that determines possible interaction between virtual content and real-world scene. Lindbauer et al. (lindlbauer2019context, ) push the AR context-awareness further by including task and cognitive load as context to adjust level of detail of AR virtual interface.
@@ 3.1. The Triad of User, Virtuality, and Reality
We began by examining the design space of AR interaction sounds. Drawing on Jain et al.’s (jain_taxonomy_2021, ) sound taxonomy in VR, we attempted a similar analysis for AR sound interaction. We first investigated several widely used code-free AR authoring tools for their AR sound capability, including Adobe Aero (adobeaero, ), Apple Reality Composer (realitycomposer, ), and Halo AR (haloar, ). We found that these tools support AR sound related to either virtual content interaction or AR-based spatial anchors, e.g. play sound when users get close to a certain position, scan a QR code, or tap on a virtual model.
The research literature in AR sound (dam2024taxonomy, ; rakkolainen2021technologies, ; filus2012using, ), however, envisions a much broader scope. E.g. Rakkolainen et al. (rakkolainen2021technologies, ) have classified AR audio into five classes: ambient, directional, musical, speech, and noise. Dam et al.’s (dam2024taxonomy, ) taxonomy of Audio Augmented Reality (AAR), defines AR-based sound across Environment Connected (sound maintaining awareness and interaction with physical environment), Goal Directed (sound assisting users’ primary goal), and Context Adapted (sound adaptive to users’ immediate reality), as well as their intersection. This taxonomy emphasizes that AR sound is often rooted in contextual information from the real world environment and user intention.
Drawing on these prior work, we introduce a novel AR sound design space that helps emphasize the sonic interplay between three key factors of AR experiences: reality, denoting the physical environment in the real world that hosts the AR experience; virtuality, referring to the virtual object(s) placed in the AR experience; and the user, representing the individual interacting with the AR experience. Below, we expand on each element of this proposed design space and indicate current levels of support in the existing code-free AR sound authoring tools (adobeaero, ; realitycomposer, ; haloar, ).
@@
Virtuality (partially supported): Virtuality refers to sounds that accompany AR events involving only virtual objects. This could be bound to an object’s change of status (e.g. a notifying sound when a virtual object shows up), from an object’s animated behavior (e.g. the mechanical noise made by a virtual dinosaur roaring), or the interaction between multiple virtual objects (e.g. two virtual balls clacking when collide). Current authoring tools support status change and animation playing as sound initiating triggers, but do not support sound for interactions between virtual objects.
User-Virtuality (well supported): User-Virtuality are the corresponding sounds that react to the users’ actions with virtual objects, e.g. a virtual dog barks when users tap on or get close to it. In all of our investigated AR creation tools, these user-initiated triggers like “Tap” and “Proximity Enter” are well supported.
User-Reality (not well supported): User-Reality refers to sounds that accompany user interactions with the physical environment via their AR devices. These sound feedbacks can improve users’ understanding of the surrounding environment. E.g. a tap sound when users tapping on a real-world table surface on their phone screen. None of our investigated AR authoring tools support user-reality actions as triggers.
Virtuality-Reality (not well supported): Virtuality-Reality denotes sound feedback that plays crucial roles in enhancing realism when virtual objects interact with the real-world environment, e.g. the crisp stomping sound of a virtual robot when it walks on a real-world glass surface. Although physics simulation between virtual models and real-world surfaces have been supported, their sound feedback remains unexplored in AR authoring tools.
User-Virtuality-Reality (not well supported): User-Virtuality-Reality are the sounds that accompany user actions involving both virtual and real elements. For example, the material-aware sliding sounds when users apply a virtual scraper on different real-world surfaces (e.g., wooden table, painted wall, or glass window). None of our investigated authoring tools can support the specification of this complex interaction.
Others Note that we excluded the domains of Reality and User from the enumeration since our interest lies in events where the AR system (e.g. smartphone or AR headset) generate the sound. Thus, sounds naturally generated by user and reality are excluded from our discussion.
Based on the above analysis and as illustrated in Fig. 2, existing AR authoring tools fail to offer functional AR sound authoring support for many of the identified dimensions, especially those involving Reality. SonifyAR aims to cover this gap.

### v3

@@ Abstract.
Sound plays a crucial role in enhancing user experience and immersiveness in Augmented Reality (AR). However, current platforms lack support for AR sound authoring due to limited interaction types, challenges in collecting and specifying context information, and difficulty in acquiring matching sound assets. We present SonifyAR, an LLM-based AR sound authoring system that generates context-aware sound effects for AR experiences. SonifyAR expands the current design space of AR sound and implements a Programming by Demonstration (PbD) pipeline to automatically collect contextual information of AR events, including virtual-content-semantics and real-world context. This context information is then processed by a large language model to acquire sound effects with Recommendation, Retrieval, Generation, and Transfer methods. To evaluate the usability and performance of our system, we conducted a user study with eight participants and created five example applications, including an AR-based science experiment, and an assistive application for low-vision AR users.
@@
conference: The ACM Symposium on User Interface Software and Technology; Oct 13–16, 2024; Pittsburgh, PAccs: Human-centered computing Interaction designccs: Human-centered computing Interactive systems and tools
Figure 1. SonifyAR is a custom AR sound authoring pipeline that generates context-matching sounds for AR events in situ using generative AI. For example, imagine sliding an AR tea cup across a real-world surface such as a wood table. SonifyAR observes this user action (a slide gesture), the action source (the user), and the action target (a virtual ceramic teacup), infers scene information such as the surface material (a wood table), and uses a custom AI backend to recommend, retrieve, generate, or sound-style transfer sound effects.
@@
SonifyAR pipeline overview. From left to right: an AR sliding operation on a virtual ceramic cup on a real-world table; A text box containing the context information of this interaction; SonifyAR use retrieve-based and generation-based methods to generate sound; The sound can be tested and experienced in AR.
@@ 1. Introduction
Sound is a critical but often overlooked element in Augmented Reality (AR). AR-based sound can improve immersion and overall user experience (Roginska and Geluso, 2017) and support depth perception and task completion (Zhou et al., 2007), search and navigation (Rumiński, 2015), and even assistance for people with low vision  (Ribeiro et al., 2012b). Despite its importance, current AR authoring platforms like Reality Composer (Apple, 2023c), Adobe Aero (Adobe, 2023a), and Unity Mars (Unity Technologies, 2023) provide only rudimentary sound support. Specifically, we identified three critical gaps with existing AR authoring tools:
@@
To address these challenges, we present SonifyAR: a context-aware AR sound authoring system using GPT-4 (OpenAI, 2024) and a text2audio diffusion model called AudioLDM (Liu et al., 2023). SonifyAR makes the following key technical advancements:
First, context collection using PbD. SonifyAR adopts a Programming by Demonstration (PbD) (Lau and Weld, 1998) pipeline to simplify the specification of complex AR interactions. The PbD pipeline enables users to demonstrate AR sound interactions while the system automatically detects the action and collects context information. For example, if a creator wants to sonify the stomping of a walking robot, they can position the virtual robot on the target (physical) surface. As the robot walks, the collision between robot’s feet and the surface, as well as the context information like the robot’s attributes and the surface’s material, is captured by SonifyAR .
Second, context as text. To utilize AR context information stemming from multiple sources (e.g. user action, virtual object, real-world environment) and in different formats (e.g. categorical, 3D shape, image), we use text as the universal medium to encompass all context information. For example, we generate an LLM prompt using the following template: “This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities]”.
@@
To author AR sound with SonifyAR, the creator initiates AR interactions and the system automatically lists sound options based on context. For example, imagine trying to make a virtual tea cup chime accordingly when being slid on a table (Figure 1). In traditional approaches, the creator would need to manually specify this sliding action and find a sound asset to match the chiming ceramic cup. But with SonifyAR, the creator could demonstrate this sliding action and the chime sounds are generated automatically for the creator to choose from.
To explore the potential of SonifyAR, we conducted two evaluations: first, a qualitative user study with eight designers to examine the usability of and reactions to the SonifyAR prototype; second, we apply SonifyAR across five key user scenarios: from AR education to improving AR headset safety. Our user study highlights the potential of SonifyAR and the generated sounds while identifying key areas of improvement such as sound quality and interface design. Our application scenarios demonstrate the breadth and potential of SonifyAR incorporating automatically acquired sounds seamlessly into various AR experiences—a practice that would otherwise take significant manual time and effort.
@@ 2.1. AR Sound Authoring
Sound in AR provides many benefits from directing the user’s attention and enhancing immersion (Roginska and Geluso, 2017; Zhou et al., 2007) to creating interactive time-varying experiences and increasing accessibility (Serafin et al., 2018; Rumiński, 2015). However, as mentioned, the current landscape of AR authoring tools reveals deficiencies in how sound is supported.
Current tools highlight a variety of AR sound authoring methods. (Monteiro et al., 2023; Nebeling and Speicher, 2018). Some—like Adobe Aero (Adobe, 2023a) and Apple Reality Composer (Apple, 2023c)—are easy to use, require no coding skills, and can only attach user-provided sound with specific AR event triggers (hal, [n. d.]; arv, [n. d.]; Adobe, 2023a; Apple, 2023c). Others—like Unity (Technologies, 2023) and Unreal Engine (Games, 2023)—are heavily code-based and require significant technical skills, but provide substantial flexibility in designing AR sound interactions, including the ability to set parameters like sound decay and code-specific conditions for activating sound assets. One common unifying approach Programming by Specification (PbS) methodology (Monteiro et al., 2023), where creators define AR sound content and play conditions during the design phase. This allows creators to precisely define the conditions for triggers and the consequent sound effects.
As mentioned in introduction, however, critical gaps exist. These gaps are partially due to the limitation of the PbS methodology, since the triggers specified in the design stage naturally lack richness in real-world context, which is crucial for AR sound realism. Also, the trade-offs between supported trigger types and the tool’s usability make most tools lack coverage for many sound-producing AR events. Additionally, all sound assets need to be manually selected or created, tasking creators with sourcing suitable sound effects to match the specific triggers. The SonifyAR system aims to address these gaps.
@@ 2.2. Sound Acquisition
Although no work has explored sound generation with AR events as input, sound generation conditioned from other input modalities like images, 3D models, text, and videos has been widely explored. Various cross-modal generative models like RNN (Zhou et al., 2018), GAN (Engel et al., 2019; Kumar et al., 2019; Ghose and Prevost, 2022), VAE (Dhariwal et al., 2020), and most recently diffusion models (Kong et al., 2020; Liu et al., 2023; Yang et al., 2023), are utilized to generate sound. Despite different model architectures, one common training goal is the mapping between latent representations of input modalities and output sound. Thus, sounds can be generated from prompts like images (Sheffer and Adi, 2023), videos (Zhou et al., 2018; Ghose and Prevost, 2022), and text (Liu et al., 2023; Yang et al., 2023). For example, FoleyGAN (Ghose and Prevost, 2022) conditions action sequences of input videos to generate visually aligned, realistic soundtracks. AudioLDM (Liu et al., 2023) utilizes the CLAP-based (Wu et al., 2023a) latent space to embed input text and generate matching sound with a diffusion model. Unlike cross-modal generative methods, another promising approach is physics-based sound synthesis (Raghuvanshi and Lin, 2006; Roodaki et al., 2017; Ren et al., 2013; Jin et al., 2020, 2022; Liu and Manocha, 2021; Diaz et al., 2022). For example, the spring-mass model (Raghuvanshi and Lin, 2006)—recently improved with deep learning (Jin et al., 2020)—can provide realistic sound simulation given a 3D model and material properties.
Besides synthesis, retrieval methods can also acquire matching sound for input conditions. For example, Koepke et al. (Koepke et al., 2022) trained cross-modal embedding models to support text-to-audio retrieval. Multi-modal models like VAST  (Chen et al., 2023) also support text-to-audio retrieval that could potentially be used to provide sound assets for AR experiences. Recent work, called Soundify (Lin et al., 2023), utilizes retrieved sound assets to sonify videos. The authors claim that although sound retrieval does not ensure coverage for all input descriptions, the retrieved sound assets usually surpass the synthesis results in terms of quality. In this case, combining generation and retrieval in the sound acquisition pipeline seems to be a natural method that integrates the merits of both. We explore this possibility with SonifyAR.
Unlike other existing sound acquisition methods, AR poses unique challenges in the matching of input condition and output sound. The multi-modal nature of AR interactions, including the information of the virtual object (e.g. a toy robot with a walking animation), real environment (e.g. a living room with a wooden table), and user action (e.g. user taps on a virtual model), needs to be carefully processed into the sound acquisition pipeline to result in matching sound assets.
@@ 2.3. Context-awareness in AR
In AR, context-awareness is an important property that allows AR content to be dynamically adjusted according to context information, such as location, scene semantics, and human factors. Context-awareness aims to adapt an application not only to the user but also to their environment and specific needs with the goal of improved usability and immersive (Krings et al., 2020). Previous research has explored various aspects of context-awareness in AR, highlighting its importance across different applications.
For example, the popular AR game Pokemón Go(Niantic, 2024) shows game contents based on user’s physical location. A more common practice in AR research is adapting virtual contents to scene semantics—essentially understanding the objects and their locations in the user’s environment. Qian et al. (Qian et al., 2022) introduced an authoring system that helps user create adaptive AR experiences to surrounding scene semantics. Lang et al. (Lang et al., 2019) analyzes scene semantics to guide the positioning of virtual agents. Similarly, Liang et al. (Liang et al., 2021) places virtual pets into AR based on scene geometry and semantics. This realm of similar work (Han et al., 2020; Tahara et al., 2020) usually captures and analyzes the surrounding environment and designs algorithms that determines possible interaction between virtual content and real-world scene. Lindbauer et al. (Lindlbauer et al., 2019) push the AR context-awareness further by including task and cognitive load as context to adjust level of detail of AR virtual interface.
@@ 3.1. The Triad of User, Virtuality, and Reality
We began by examining the design space of AR interaction sounds. Drawing on Jain et al.’s (Jain et al., 2021) sound taxonomy in VR, we attempted a similar analysis for AR sound interaction. We first investigated several widely used code-free AR authoring tools for their AR sound capability, including Adobe Aero (Adobe, 2023a), Apple Reality Composer (Apple, 2023c), and Halo AR (hal, [n. d.]). We found that these tools support AR sound related to either virtual content interaction or AR-based spatial anchors (e.g.,play a sound when a user gets close to a certain position, scans a QR code, or taps on a virtual model).
The AR sound research literature (Dam et al., 2024; Rakkolainen et al., 2021; Filus and Rambli, 2012), however, envisions a much broader scope. Rakkolainen et al. (Rakkolainen et al., 2021) classified AR audio into five classes: ambient, directional, musical, speech, and noise. Dam et al.’s (Dam et al., 2024) taxonomy of Audio Augmented Reality (AAR), defines AR-based sound across Environment Connected (sound maintaining awareness and interaction with physical environment), Goal Directed (sound assisting users’ primary goal), and Context Adapted (sound adaptive to users’ immediate reality), as well as their intersection. This taxonomy emphasizes that AR sound is often rooted in contextual information from the real world environment and user intention.
Drawing on this prior work, we introduce a novel AR sound design space that helps emphasize the sonic interplay between three key factors of AR experiences: reality, denoting the physical environment in the real world that hosts the AR experience; virtuality, referring to the virtual object(s) placed in the AR experience; and the user, representing the individual interacting with the AR experience. Below, we expand on each element of this proposed design space and indicate current levels of support in the existing code-free AR sound authoring tools (Adobe, 2023a; Apple, 2023c; hal, [n. d.]).
@@
[…]
Sound design space shown as a hexagon between User, Virtuality and Reality. SonifyAR can support all listed sound design cases, while the existing work can only support the User-Virtuality and Virtuality cases.
Virtuality (partially supported): Virtuality refers to sounds that accompany AR events involving only virtual objects. This could be bound to an object’s change of status (e.g., a notifying sound when a virtual object shows up), from an object’s animated behavior (e.g., the mechanical noise made by a virtual dinosaur roaring), or the interaction between multiple virtual objects (e.g., two virtual balls clacking when they collide). Current authoring tools support status changes and animation playing as sound-initiating triggers, but do not support sound for interactions between virtual objects.
User-Virtuality (well supported): User-Virtuality are the corresponding sounds that react to the users’ actions with virtual objects (e.g., a virtual dog barks when users tap on or gets close to it). In all of our investigated AR creation tools, these user-initiated triggers like “Tap” and “Proximity Enter” are well supported.
User-Reality (not well supported): User-Reality refers to sounds that accompany user interactions with the physical environment via their AR devices. This sound feedback can improve users’ understanding of the surrounding environment (e.g., an appropriate tap sound when users tap on a real-world table surface via their phone screen). None of the AR authoring tools we investigated support user-reality actions as triggers.
Virtuality-Reality (not well supported): Virtuality-Reality denotes sound feedback that plays a crucial role in enhancing realism when virtual objects interact with the real-world environment (e.g., the crisp stomping sound of a virtual robot when it walks on a real-world glass surface). Although physics simulations between virtual models and real-world surfaces have been supported, their sound feedback remains unexplored in AR authoring tools.
User-Virtuality-Reality (not well supported): User-Virtuality-Reality are the sounds that accompany user actions involving both virtual and real elements. For example, the material-aware sliding sounds when users apply a virtual scraper on different real-world surfaces (e.g., a wooden table, painted wall, or glass window). None of our investigated authoring tools can support the specification of this complex interaction.
Others: We exclude the domains of Reality and User from the enumeration since our interest lies in events where the AR system (e.g. smartphone or AR headset) generate the sound. Thus, sounds naturally generated by user and reality (e.g., a user physically knocking on a wooden table) are excluded from our discussion.
Based on the above analysis and as illustrated in Fig. 2, existing AR authoring tools fail to offer functional AR sound authoring support for many of the identified dimensions, especially those involving Reality. SonifyAR aims to address this gap.